MTH 1001-Quantitative Reasoning- Test 4 Use the information in the given chart for question # 20 fractions You may leave your answers as unreduced Marital Status of the U.S. Population, ages 15 or older, 2010, in millions Married Widowed Total Never Married 40 34 74 Divorced Male Female Total 65 65 130 10 14 24 118 124 242 20) If one person is randomly selected from the population described in the table, find: a) P(person is divorced) = b) P(person is female or married)- e) P(person is not widowed)- d) P(person is married, given they are female)- e) P(person is male, given they are divorced)-
Answer #1

a) P(divorced) = 24 / 242

b) P(female or married) = P(female) + P(married) - P(female and married)

                                      = 124/242 + 130/242 - 65/242

                                      = 189/242

c) P(not widowed) = 1 - P(widowed) = 1 - 14/242 = 228/242

d) P(married | female) = P(married and female) / P(female)

                                    = (65/242) / (124/242)

                                    = 65/124

e) P(male | divorced) = P(male and divorced) / P(divorced)

                                  = (10/242) / (24/242)

                                  = 10/24